Sometimes I buy M&M’s instead of peanuts (like I did in the example yesterday). Find the

SAMPLE standard deviation of the number of chocolates in these six bags. You are more than welcome to create a chart below the data to help you find the solution.

21, 30, 25, 19, 24, 28


1
Expert's answer
2021-01-14T20:01:40-0500

Here the given data is as follows "21,30,25,19,24,28"

Number of observation "N=6"

Mean of the above data "\mu =\frac{\Sigma X_i}{N}=\frac{(21+30+25+19+24+28)}{6}=24.5"

Standard deviation of data "\sigma =\sqrt{\frac{\Sigma (X_i-\mu)^2}{N}}"

For this we draw a table below




Therefore Standard Deviation("\sigma" ) "=\sqrt {\frac {(12.25+30.25+0.25+30.25+0.25+12.25)}{6}}"

"=\sqrt{\frac{85.5}{6}}=\sqrt{14.25}" "\approx 3.77" (correct upto two decimal places)

Hence required Standard Deviation ("\sigma") "=3.77" (approximately).